A first order reaction is carried out with an initial concentration of 10 mol per litre and 80% of the reactant changes into the product. Now if the same reaction is carried out with an initial concentration of 5 mol/L for the same period the % of th reactant changing to the product is-
a.40   b.80    c.160

Grade:

1 Answers

Arun
5 years ago
Dear Tanusri
 
 
For the 10ML-1 concentration,
k = ln(10/2)/t.....(1)
and
for 5ML-1
k = ln(5/x)/t..............(2)
From both the eqns. we get,
5 = 5/x
x = 1
that means % that gets converted to product, 
4*100/5 => 80%
 
 
Regards
Arun (askIITians forum expert)